OpenStudy (anonymous):

5962*2568/457*9632-10?

OpenStudy (kkrpbra):

that dont make sence

OpenStudy (anonymous):

yes an thats not how you spell since!!

OpenStudy (anonymous):

makes perfect sense PEMDAS

OpenStudy (anonymous):

lol do u know the answer?

OpenStudy (anonymous):

got a calculator?

OpenStudy (anonymous):

Can we get clarification on numerator and denominator?

OpenStudy (anonymous):

i got ! 3.478206030513786

OpenStudy (anonymous):

............

OpenStudy (anonymous):

Assuming from the lack of parentheses that it means \[5962 \times 2568 \div 547 \times 9632 - 10\] and not \[5962 \times \frac{2568}{547} \times 9632 - 10\] Then the order of operations is \[\frac{5962 \times 2568}{547} \space \times 9632 - 10\]

OpenStudy (anonymous):

ITS THE SAME THINGGGGG

OpenStudy (anonymous):

i believe its called the communative property of multiplication

OpenStudy (anonymous):

Oops, meant for that second one to be \[\frac{5962 \times 2568}{547 \times 9632} \space -10\] Fumbled with the LaTeX..

OpenStudy (kkrpbra):

they really trying 2 answer this smh

OpenStudy (anonymous):

really its pemdas SO WORK IT OUT!!!!